Let the "pdf" of X be given by "f(x)=0.25exp(-0.25x),x>0"

Show that "E(X)" =4 and "Var(X)" =16.


1
Expert's answer
2021-05-07T10:30:31-0400

We have given that,


"f(x) = 0.25e^{-0.25x},x>0"


We have to find the value of E(X),


We know,


"E(X) = \int_{0}^{\infin} x0.25e^{-0.25x}dx"


"= 0.25\int_{0}^{\infin} xe^{-0.25x} dx"


"= 0.25[-4xe^{-0.25x}-16e^{-0.25x}]_{0}^{\infin}"


"= 0.25 \times 16"


"= 4"


Hence, "E(X) = 4"


"Var(X) = E(X^2)-{E(X)}^2"


"E(X^2) = 0.25 \int_{0}^{\infin} x^2e^{-0.25x}d" x


"= 0.25[-4x^2e^{-0.25x}-32xe^{-0.25x}-128e^{-0.25x}]_{0}^{\infin}"


"= 0.25 \times 128"


"= 32"


Hence, "Var(X) = 32-16 = 16"